# The stochastic discrete nonlinear Schrödinger equation: microscopic derivation and finite-temperature phase transition

### v1

@@ III.2 Localised and disordered phase
Figure 5: Space-time density plots of the pattern of amplitudes, $|c_{n}(t)|^{2}$, in the stationary state as obtained from numerical simulations of eq. (11) with stepsize $\tau=0.01$ for a system of size $L=64$, noise strength $\sigma=0.3$ and $\alpha=16$. Temperature values: (a) $\bar{\beta}_{low}=2.175\quad(\beta_{low}=17.4)$, (b) $\bar{\beta}_{crit}=2.4\quad(\beta_{crit}=19.2)$, and (c) $\bar{\beta}_{high}=2.5\quad(\beta_{high}=20)$(see as well Figure 5(a)), Colour coding: from blue ($|c|^{2}=0.01$) to yellow ($|c|^{2}=0.99$). Panels (d), (e), and (f) show the corresponding time evolution of the total energy $E(t)$. Data have been obtained using a localised initial condition $c_{n}(0)=\delta_{n,32}$, skipping a transient of length about $\Delta t=1.6\times 10^{5}$ to ensure stationary behaviour.
@@
For a quantitative characterisation of patterns and phases, we resort to a computation of the distribution of amplitudes, $W(|c|^{2})$. The data in Figure 6(a) show an exponentially decaying distribution for negative and sufficiently small positive $\bar{\beta}$ values, with a peak at large values of $|c|^{2}$ building up as the critical region is approached and becoming dominant in the high-$\beta$ phase. These results of the stochastic model are completely in line with results reported for the Hamiltonian dynamics of the DNSE early on in the context of negative temperature states [rasmussen_statistical_2000, rumpf_transition_2008]. Thus, the features of the stochastic motion cover important aspects of the deterministic model. A basic characterisation of the temporal properties of the stationary behaviour is provided through the energy autocorrelation function. In the low- and the high-$\beta$ phase energy correlations decay exponentially on a fast time scale (see Figure 6(b)), while, as expected, correlations slow down considerably in the critical region (cf. as well the time traces in Figure 5, and Figure 2(b) for the dependence of the correlation function on the noise strength $\sigma$ in the disordered phase).
@@
Figure 6: (a) Probability distribution $W(|c|^{2})$ of amplitudes in the stationary state, obtained for a system of size $L=64$, noise strength $\sigma=0.3$, $\alpha=16$, and different values of the temperature, including the three values used in Figure 5(a-c). Data have been computed from a time series of length about $T=2.6\times 10^{6}$ with sampling rate $\tau=0.01$, skipping a transient of duration about $\Delta t=1.6\times 10^{5}$ to ensure stationary behaviour. (b) Autocorrelation function of the energy for the parameter setting of part (a), showing that the correlation time is longest in the middle of the phase transition region. Data have been obtained from time series of length about $T=2.6\times 10^{6}$ skipping transients of length about $\Delta t=1.6\times 10^{5}$.
@@
So far, we have focused on the stationary behaviour of the different phases. A key aspect of nonequilibrium dynamics consists in the formation of localised structures in the localised phase starting from a random initial condition, as well as the decay of localised structures in the disordered phase when starting from a localised initial condition (cf. e.g. [Bray_AP93] for a comprehensive overview of these aspects in the context of condensed matter physics). A simple quantity that captures such an aspect is the so-called inverse participation ratio $Y(t)=\sum_{n=1}^{L}|c_{n}(t)|^{4}$. Small values indicate a delocalised structure, while the emergence of a localised structure is signalled by inverse participation ratios approaching the maximum value of one. Figure 7 shows the time evolution of the inverse participation ratio in the two scenarios mentioned above. If we start with high $\beta$ values, that means in the localised phase, from random initial conditions drawn from the distribution $\delta(N-1)$, the inverse participation ratio $Y(t)$ increases with time, indicating that localised structures appear when approaching the stationary state, Figure 7(a). One observes a non-monotonic dependence of the process on the noise strength $\sigma$, and there seems to appear an optimal noise strength which supports the formation of localised structures. At first sight, such a kind of phenomenon reminiscent of noise-induced formation of structures [HoLe:84] seems plausible. However, looking at the low-$\beta$ case, i.e. the random phase, and starting from a localised initial condition $c_{n}(0)=\delta_{n,32}$ we see the decay of this structure, signalled by a decay of the inverse participation ratio, see Figure 7(b). Again, it seems that the same optimal noise strength speeds up this process. Unlike the simple diffusive picture mentioned in section II.2 the impact of noise is more sophisticated in this setup and comes from an interaction of the deterministic nonlinear dynamics with the stochastic force in eq. (11), which so far defies a simple heuristic reasoning, let alone a proper formal explanation.
@@ IV Discussion and Conclusion
The gist of the phase transition itself can be captured by a simple mean-field model that gives an approximate expression of the canonical partition function $Z$ in dependence on a suitable parameter. Minimization of $F=-k_{B}T\ln Z$ with respect to this parameter then yields the thermodynamically stable phases. We assume that $\alpha$ is large enough so that in the low-temperature phase the breather can be approximated by a state where, on average, the amplitude $|c_{n}|^{2}$ on one site is larger than on the others. We implement this by setting $\langle|c_{1}|^{2}\rangle=a$ and $\langle|c_{n}|^{2}\rangle=b$ for $n>1$, where normalisation requires $\langle N\rangle=a+(L-1)b=1$. This ansatz accommodates both phases as the high-temperature phase is characterised by $a=b=1/L$. The low-temperature phase breaks translational symmetry since the mean amplitude at one site differs from that of the others. The difference aba-b can be interpreted as the order parameter associated with the phase transition. The normalisation condition constrains the parameter $b$ to the range $[0,1/L]$ to leading order in system size, as $a\geq 0$ and $b\geq 0$. The energy of the system can be estimated by
@@
keeping only leading-order terms in $L$ and using the abbreviation $x=bL\in[0,1]$. In writing down eq. (12) we neglected the nearest-neighbour coupling terms. To obtain the partition integral $Z$, we estimate the degeneracy of each energy value by making the plausible assumption that the range of the amplitudes $|c_{n}|^{2}$ is of the same order as their mean, which is $b$. We thus obtain the estimate
@@
The scaled free energy, eq. (15) has a global minimum at $x=1$ for $\bar{\beta}<2$ and develops a second local minimum at $\bar{\beta}=2$ which becomes the global minimum for $\bar{\beta}>\bar{\beta}_{c}\simeq 2.46$, see Figure 8. The value of $x=Lb$ that minimises $F$ jumps from $1$ to the vicinity of $0.3$ at the critical temperature determined by $\bar{\beta}_{c}$. The jump of $x$ (and therefore the jump of the order parameter aba-b) means that the phase transition is of first order. The inner minimum of the free energy (15) becomes metastable for $2<\bar{\beta}<\bar{\beta}_{c}$ but persists down to $\bar{\beta}=2$, where this minimum vanishes. The outer minimum at 1 remains metastable for $\beta>\bar{\beta}_{c}$. The energy associated with the global minimum $x_{min}$ is obtained from eq. (12) and shown above in Figure 4 as a black dash-dotted line. The location $\bar{\beta}_{c}$ of the jump agrees pretty well with the simulation data in this figure, and so does the entire low-temperature phase, $\bar{\beta}>\bar{\beta}_{c}$, including the metastable region.
Figure 8: (a) Dependence of the nontrivial part of the free energy, eq. (15), on the order parameter $x$, for different values of the scaled temperature $\bar{\beta}$. (b) Global (solid) and local (dash-dotted) minima $x_{min}$ of the free energy, eq. (15), in dependence of the scaled temperature $\bar{\beta}$. The dashed vertical line indicates the critical value $\bar{\beta}_{c}$.
@@
The instability of the homogeneous state of the system over a large energy interval might help to understand results obtained in microcanonical simulations, where the energy is constrained to have the same value at all times. According to our mean-field theory, such energy values can be realised by an $x=bL$ value that does not minimise the free energy $F$. This state is already a symmetry-broken state since the mean amplitude at one site is larger than that of the other sites. In the canonical ensemble, this solution becomes unstable and is replaced by a phase coexistence, a situation that bears some resemblance to the Maxwell construction performed for the solution of the van der Waals equation in the parameter range where the homogeneous solution is unstable.
@@
For energies $E>E_{c}(\alpha)$ the DNSE is in the disordered phase whereas for $E<E_{c}(\alpha)$, the DNSE is either in the phase coexistence region or in the localised phase. This critical line $E_{c}(\alpha)$ is shown in the phase diagram Figure 9 together with its conjugate counterpart for $\alpha<0$. This theoretical prediction agrees quite well with results obtained from numerical simulations. The data points next to this line have been obtained from simulations in a system of size $L=1024$ with noise strength $\sigma=0.3$ skipping transients of duration $\Delta t=3.4\times 10^{5}$ and computing energy averages over a time interval of length $T=3.4\times 10^{5}$ (to cope with stiffness due to large $\beta$ values the midpoint step in the numerical integration procedure, see appendix D, has been occasionally replaced by more stable schemes). The dots indicate the energy the system jumps to when the low-temperature phase in numerical simulations becomes unstable in an adiabatic parameter downsweep.
@@
Our stochastic equation of motion, eq. (11), includes the noise strength $\sigma$ as an additional parameter. the stochastic model, eq. (11). The non-monotonic dependence of the transient behaviour on $\sigma$, see Figure 7, resulting in an optimal noise strength, is an unexpected property. Such a feature might be caused by a nontrivial interaction between the noise and internal time scales of the deterministic dynamics, reminiscent of stochastic resonance. A proper explanation, which is beyond the scope of the present paper, may shed some light on the intricate nonequilibrium properties of the DNSE.
@@ Appendix A Microscopic derivation
where $q$ denotes the wavenumers, $b_{q}$ the phonon amplitudes, and $\omega_{q}$ the dispersion relation. Overall, the dynamics should still preserve the normalisation $N$, which puts a constraint on the coupling between the subsystem and the heat bath. Here we chose the lowest-order term that is compatible with the symmetries and conservation laws of the system, which gives a coupling that is linear in the phonon amplitudes $b_{q}$ and linear in the amplitudes $|c_{n}|^{2}$ (to preserve the normalisation $N$),
@@
$\{F,G\}$ $=$ $\sum_{n}i\Big(\frac{\partial F}{\partial\bar{c}_{n}}\frac{\partial G}{\partial c_{n}}-\frac{\partial F}{\partial c_{n}}\frac{\partial G}{\partial\bar{c}_{n}}\Big)$ (20)
$+\sum_{q}i\Big(\frac{\partial F}{\partial\bar{b}_{q}}\frac{\partial G}{\partial b_{q}}-\frac{\partial F}{\partial b_{q}}\frac{\partial G}{\partial\bar{b}_{q}}\Big)\,.$
@@
$\gamma_{n-m}(t-s)=i\sum_{q}|g_{q}|^{2}\Big[e^{-i(n-m)q}e^{i\omega_{q}(t-s)}$
$\qquad\qquad-e^{\,i(n-m)q}e^{-i\omega_{q}(t-s)}\Big].$ (24)
@@
Comparing the expression (26) for the noise correlation function with expression (A) for the memory kernel, we obtain the relation
@@
This equation constitutes a fluctuation-dissipation theorem (or a generalised Einstein relation) for our model. It ensures that noise and damping are consistent with thermodynamic equilibrium. Substituting relation (27) in the right hand side of eq. (A) gives, using an integration by parts and considering times large compared to $\tau_{c}$
@@
$\dot{c}_{n}(t)$ $=$ $i\frac{\partial H_{S}}{\partial\bar{c}_{n}}+ic_{n}(t)f_{n}(t)$
$+i\beta\frac{\sigma^{2}}{2}c_{n}(t)\left(\dot{c}_{n}(t)\bar{c}_{n}(t)+c_{n}(t)\dot{\bar{c}}_{n}(t)\right)$
@@
Here, we have assumed, for simplicity, that any bath-induced interaction terms caused by $\Gamma_{n-m}(s)$ with $n\neq m$ can be neglected, which is an implicit assumption about the coupling coefficients $g_{q}$. The last term in eq. (A) which represents the dissipation in the system is a priori a Gilbert-type damping term (see e.g. [Kiku_JAP56] and [ElFi_PA83], or [HiMO_PRL09] for a recent discussion of the issue). We may solve eq. (A) for the time derivatives if we cast that equation in the form
$\left(\begin{array}[]{cc}\displaystyle 1-i\beta\frac{\sigma^{2}}{2}|c_{n}(t)|^{2}&-i\beta\frac{\sigma^{2}}{2}c_{n}^{2}(t)\\ \displaystyle i\beta\frac{\sigma^{2}}{2}\bar{c}_{n}^{2}(t)&1+i\beta\frac{\sigma^{2}}{2}|c_{n}(t)|^{2}\end{array}\right)\left(\begin{array}[]{c}\dot{c}_{n}(t)\\ \dot{\bar{c}}_{n}(t)\end{array}\right)$ (38)
$=$ $\left(\begin{array}[]{c}\displaystyle i\frac{\partial H_{S}}{\partial\bar{c}_{n}}+i\sigma c_{n}(t)\xi_{n}(t)\\ \displaystyle-i\frac{\partial H_{S}}{\partial c_{n}}-i\sigma\bar{c}_{n}(t)\xi_{n}(t)\end{array}\right)$ (41)

### v2

@@ 3.2 Localised and disordered phase
Figure 5: Space-time density plots of the pattern of weights, $|c_{n}(t)|^{2}$, in the stationary state as obtained from numerical simulations of eq. (11) with stepsize $\tau=0.01$ for a system of size $L=64$, noise strength $\sigma=0.3$ and $\alpha=16$. Scaled inverse temperature values: (a) $\bar{\beta}_{low}=2.175\quad(\beta_{low}=17.4)$, (b) $\bar{\beta}_{crit}=2.4\quad(\beta_{crit}=19.2)$, and (c) $\bar{\beta}_{high}=2.5\quad(\beta_{high}=20)$(see as well Figure 5(a)), Colour coding: from blue ($|c|^{2}=0.01$) to yellow ($|c|^{2}=0.99$). Panels (d), (e), and (f) show the corresponding time evolution of the total energy $E(t)$. Data have been obtained using a localised initial condition $c_{n}(0)=\delta_{n,32}$, skipping a transient of length about $\Delta t=1.6\times 10^{5}$ to ensure stationary behaviour.
@@
For a quantitative characterisation of patterns and phases, we resort to a computation of the distribution of weights, $W(|c|^{2})$. The data in Figure 6(a) show an exponentially decaying distribution for negative and sufficiently small positive $\bar{\beta}$ values, with a peak at large values of $|c|^{2}$ building up as the critical region is approached and becoming dominant in the high-$\beta$ phase. These results of the stochastic model are completely in line with results reported for the Hamiltonian dynamics of the DNSE early on in the context of negative temperature states [rasmussen_statistical_2000, rumpf_transition_2008]. Thus, the features of the stochastic motion cover important aspects of the deterministic model. A basic characterisation of the temporal properties of the stationary behaviour is provided through the energy autocorrelation function. In the low- and the high-$\beta$ phase energy correlations decay exponentially on a fast time scale (see Figure 6(b)), while, as expected, correlations slow down considerably in the critical region (cf. as well the time traces in Figure 5, and Figure 2(b) for the dependence of the correlation function on the noise strength $\sigma$ in the disordered phase).
Refer to captionRefer to caption
@@
Figure 6: (a) Probability distribution $W(|c|^{2})$ of weights in the stationary state, obtained for a system of size $L=64$, noise strength $\sigma=0.3$, $\alpha=16$, and different values of the temperature, including the three values used in Figure 5(a-c). Data have been computed from a time series of length about $T_{sim}=2.6\times 10^{6}$ with sampling rate $\tau=0.01$, skipping a transient of duration about $\Delta t=1.6\times 10^{5}$ to ensure stationary behaviour. (b) Autocorrelation function of the energy for the parameter setting of part (a), showing that the correlation time is longest in the middle of the phase transition region. Data have been obtained from time series of length about $T_{sim}=2.6\times 10^{6}$ skipping transients of length about $\Delta t=1.6\times 10^{5}$.
@@
So far, we have focused on the stationary behaviour of the different phases. A key aspect of nonequilibrium dynamics consists in the formation of localised structures in the localised phase starting from a random initial condition, as well as the decay of localised structures in the disordered phase when starting from a localised initial condition (cf. e.g. [Bray_AP93] for a comprehensive overview of these aspects in the context of condensed matter physics). A simple quantity that captures such an aspect is the so-called inverse participation ratio $Y(t)=\sum_{n=1}^{L}|c_{n}(t)|^{4}$. Small values indicate a delocalised structure, while the emergence of a localised structure is signalled by inverse participation ratios approaching the maximum value of one. Figure 7 shows the time evolution of the inverse participation ratio in the two scenarios mentioned above. If we start with high $\beta$ values, that means in the localised phase, from random initial conditions drawn from the distribution $\delta(N-1)$, the inverse participation ratio $Y(t)$ increases with time, indicating that localised structures appear when approaching the stationary state, Figure 7(a). One observes a non-monotonic dependence of the process on the noise strength $\sigma$, and there seems to appear an optimal noise strength which supports the formation of localised structures. At first sight, such a kind of phenomenon reminiscent of noise-induced formation of structures [HoLe:84] seems plausible. However, looking at the low-$\beta$ case, i.e. the random phase, and starting from a localised initial condition $c_{n}(0)=\delta_{n,32}$ we see the decay of this structure, signalled by a decay of the inverse participation ratio, see Figure 7(b). Again, it seems that the same optimal noise strength speeds up this process. Unlike the simple diffusive picture mentioned in section 2.2 the impact of noise is more sophisticated in this setup and comes from an interaction of the deterministic nonlinear dynamics with the stochastic force in eq. (11), which so far defies a simple heuristic reasoning, let alone a proper formal explanation.
Refer to captionRefer to caption
@@ 4 Discussion and Conclusion
The gist of the phase transition itself can be captured by a simple mean-field model that gives an approximate expression of the canonical partition function $Z$ in dependence on a suitable parameter. Minimization of $F=-k_{B}T\ln Z$ with respect to this parameter then yields the thermodynamically stable phases. We assume that $\alpha$ is large enough so that in the low-temperature phase the breather can be approximated by a state where, on average, the weight $|c_{n}|^{2}$ on one site is larger than on the others. We implement this by setting $\langle|c_{1}|^{2}\rangle=a$ and $\langle|c_{n}|^{2}\rangle=b$ for $n>1$, where normalisation requires $\langle N\rangle=a+(L-1)b=1$. This ansatz accommodates both phases as the high-temperature phase is characterised by $a=b=1/L$. The low-temperature phase breaks translational symmetry since the mean weight at one site differs from that of the others. The difference aba-b can be interpreted as the order parameter associated with the phase transition. The normalisation condition constrains the parameter $b$ to the range $[0,1/L]$ to leading order in system size, as $a\geq 0$ and $b\geq 0$. The energy of the system can be estimated by
@@
keeping only leading-order terms in $L$ and using the abbreviation $x=bL\in[0,1]$. In writing down eq. (12) we neglected the nearest-neighbour coupling terms. To obtain the partition integral $Z$, we estimate the degeneracy of each energy value by making the plausible assumption that the range of the weights $|c_{n}|^{2}$ is of the same order as their mean, which is $b$. We thus obtain the estimate
@@
The scaled free energy, eq. (15) has a global minimum at $x=1$ for $\bar{\beta}<2$ and develops a second local minimum at $\bar{\beta}=2$ which becomes the global minimum for $\bar{\beta}>\bar{\beta}_{c}\simeq 2.46$, see Figure 8. The value of $x=Lb$ that minimises $F$ jumps from $1$ to the vicinity of $0.3$ at the critical temperature determined by $\bar{\beta}_{c}$. This means that approximately 70 percent of all weight is concentrated in one site. The jump of $x$ (and therefore the jump of the order parameter aba-b) means that the phase transition is of first order. The inner minimum of the free energy (15) becomes metastable for $2<\bar{\beta}<\bar{\beta}_{c}$ but persists down to $\bar{\beta}=2$, where this minimum vanishes. The outer minimum at 1 remains metastable for $\beta>\bar{\beta}_{c}$. The energy associated with the global minimum $x_{min}$ is obtained from eq. (12) and shown above in Figure 4(b) as a black dash-dotted line. The location $\bar{\beta}_{c}$ of the jump agrees pretty well with the simulation data in this figure, and so does the entire low-temperature phase, $\bar{\beta}>\bar{\beta}_{c}$, including the metastable region.
Refer to captionRefer to caption
Figure 8: (a) Dependence of the nontrivial part of the free energy, eq. (15), on the order parameter $x$, for different values of the scaled inverse temperature $\bar{\beta}$. (b) Global (solid) and local (dash-dotted) minima $x_{min}$ of the free energy, eq. (15), in dependence of the scaled inverse temperature $\bar{\beta}$. The dashed vertical line indicates the critical value $\bar{\beta}_{c}$.
@@
Interestingly, our results for the order parameter $x_{min}$, for the critical inverse scaled temperature $\bar{\beta}_{c}$, and for the free energy function $\Phi(x)$ evaluated at $x_{min}$ agree (apart from an additive constant in the free energy) with the results of an exact analytical calculation for the full DNSE (including nearest-neighbor coupling) performed by Chatterjee and Kirkpatrick [chatterjee2012probabilistic] in the limit where the number of lattice sites goes to infinity at fixed average weight per site (corresponding in our notation to a fixed $\alpha/L$). This means that our simple mean-field theory becomes exact in this limit.
@@
The instability of the homogeneous state of the system over a large energy interval might help to understand results obtained in microcanonical simulations, where the energy is constrained to have the same value at all times. According to our mean-field theory, such energy values can be realised by an $x=bL$ value that does not minimise the free energy $F$. This state is already a symmetry-broken state since the mean weight at one site is larger than that of the other sites. In the canonical ensemble, this solution becomes unstable and is replaced by a phase coexistence, a situation that bears some resemblance to the Maxwell construction performed for the solution of the van der Waals equation in the parameter range where the homogeneous solution is unstable.
@@
For energies $E>E_{c}(\alpha)$ the DNSE is in the disordered phase whereas for $E<E_{c}(\alpha)$, the DNSE is either in the phase coexistence region or in the localised phase. This critical line $E_{c}(\alpha)$ is shown in the phase diagram Figure 9 together with its conjugate counterpart for $\alpha<0$. This theoretical prediction agrees quite well with results obtained from numerical simulations. The data points next to this line have been obtained from simulations in a system of size $L=1024$ with noise strength $\sigma=0.3$ skipping transients of duration $\Delta t=3.4\times 10^{5}$ and computing energy averages over a time interval of length $T_{sim}=3.4\times 10^{5}$ (to cope with stiffness due to large $\beta$ values the midpoint step in the numerical integration procedure, see appendix D, has been occasionally replaced by more stable schemes). The dots indicate the energy the system jumps to when the low-temperature phase in numerical simulations becomes unstable in an adiabatic parameter downsweep.
@@
Our stochastic equation of motion, eq. (11), includes the noise strength $\sigma$ as an additional parameter. The non-monotonic dependence of the transient behaviour on $\sigma$, see Figure 7, resulting in an optimal noise strength, is an unexpected property. Such a feature might be caused by a nontrivial interaction between the noise and internal time scales of the deterministic dynamics, reminiscent of stochastic resonance. Stochastic resonance is normally triggered by a resonance phenomenon between two time scales, often caused by a deterministic external drive and an internal stochastic motion which depends on noise strength, resulting in an enhanced periodic signature in output signals occurring at on optimal noise strength. In our case we observe a non-monotonic dependence of transient times on the noise strength. Transient times become minimal at an optimal noise strength, a feature which is certainly linked with the internal dynamics of the DNSE. Having said that there is no obvious resonance phenomenon between two different time scales visible. A proper explanation, which is beyond the scope of the present paper, may shed some light on the intricate nonequilibrium properties of the DNSE.
@@
[…]
No new data were created or analysed in this study.
[…]
None.
@@ Appendix A Microscopic derivation
where $q$ denotes the wavenumers, $b_{q}$ the phonon amplitudes, and $\omega_{q}$ the dispersion relation. Overall, the dynamics should still preserve the normalisation $N$, which puts a constraint on the coupling between the subsystem and the heat bath. Here we chose the lowest-order term that is compatible with the symmetries and conservation laws of the system, which gives a coupling that is linear in the phonon amplitudes $b_{q}$ and linear in the weights $|c_{n}|^{2}$ (to preserve the normalisation $N$),
@@
$\{F,G\}$ $=$ $\sum_{n}i\Big(\frac{\partial F}{\partial\bar{c}_{n}}\frac{\partial G}{\partial c_{n}}-\frac{\partial F}{\partial c_{n}}\frac{\partial G}{\partial\bar{c}_{n}}\Big)+\sum_{q}i\Big(\frac{\partial F}{\partial\bar{b}_{q}}\frac{\partial G}{\partial b_{q}}-\frac{\partial F}{\partial b_{q}}\frac{\partial G}{\partial\bar{b}_{q}}\Big)\,.$ (20)
@@
$\gamma_{n-m}(t-s)=i\sum_{q}|g_{q}|^{2}\Big[e^{-i(n-m)q}e^{i\omega_{q}(t-s)}-e^{\,i(n-m)q}e^{-i\omega_{q}(t-s)}\Big].$ (24)
@@
Comparing the expression (26) for the noise correlation function with expression (24) for the memory kernel, we obtain the relation
@@
This equation constitutes a fluctuation-dissipation theorem (or a generalised Einstein relation) for our model. It ensures that noise and damping are consistent with thermodynamic equilibrium. Substituting relation (27) in the right hand side of eq. (22) gives, using an integration by parts and considering times large compared to $\tau_{c}$
@@
$\dot{c}_{n}(t)$ $=$ $i\frac{\partial H_{S}}{\partial\bar{c}_{n}}+ic_{n}(t)f_{n}(t)+i\beta\frac{\sigma^{2}}{2}c_{n}(t)\left(\dot{c}_{n}(t)\bar{c}_{n}(t)+c_{n}(t)\dot{\bar{c}}_{n}(t)\right)$ (31)
@@
Here, we have assumed, for simplicity, that any bath-induced interaction terms caused by $\Gamma_{n-m}(s)$ with $n\neq m$ can be neglected, which is an implicit assumption about the coupling coefficients $g_{q}$. The last term in eq. (31) which represents the dissipation in the system is a priori a Gilbert-type damping term (see e.g. [Kiku_JAP56] and [ElFi_PA83], or [HiMO_PRL09] for a recent discussion of the issue). We may solve eq. (31) for the time derivatives if we cast that equation in the form
$\left(\begin{array}[]{cc}\displaystyle 1-i\beta\frac{\sigma^{2}}{2}|c_{n}(t)|^{2}&-i\beta\frac{\sigma^{2}}{2}c_{n}^{2}(t)\\ \displaystyle i\beta\frac{\sigma^{2}}{2}\bar{c}_{n}^{2}(t)&1+i\beta\frac{\sigma^{2}}{2}|c_{n}(t)|^{2}\end{array}\right)\left(\begin{array}[]{c}\dot{c}_{n}(t)\\ \dot{\bar{c}}_{n}(t)\end{array}\right)=\left(\begin{array}[]{c}\displaystyle i\frac{\partial H_{S}}{\partial\bar{c}_{n}}+i\sigma c_{n}(t)\xi_{n}(t)\\ \displaystyle-i\frac{\partial H_{S}}{\partial c_{n}}-i\sigma\bar{c}_{n}(t)\xi_{n}(t)\end{array}\right)$ (40)